OpenStudy (jkbo):

Which equation is in standard form and represents a line with slope –1 through the point (–3, 5)?

OpenStudy (jkbo):

A. x+y=2 B. y=-x+5 c. y=-x+2 D.-x+y=5

OpenStudy (anonymous):

You would start with the point-slope form of the equation for a line and then go to the standard form. Point-slope form: y - y1 = m(x - x1) where you have a given point of (-3, 5) which is your (x1, y1) and "m" is your slope. y - 5 = (-1)[x - (-3)] -> y - 5 = -x - 3 Now, get the "x" and "y" terms over to the left by themselves. x + y = 2

OpenStudy (anonymous):

Are you there, @jkbo ?

OpenStudy (jkbo):

yes thank you